OpenStudy (anonymous):

square root of -6 mutliplied by the square root of -384

OpenStudy (gorv):

\[\sqrt{-6}*\sqrt{-384}\]

OpenStudy (gorv):

\[\sqrt{-6*(-384)}=\sqrt{6*384}=\sqrt{2304}\]

OpenStudy (gorv):

when two minus sign are multiplied they become plus

OpenStudy (anonymous):

\[\sqrt{-6} \sqrt{-384}\]

OpenStudy (anonymous):

that is the equation

OpenStudy (gorv):

when two number have same power then can be return together

OpenStudy (gorv):

x^2*y^2=(x*y)^2

OpenStudy (gorv):

similarly for square root ..dont worry its correct

OpenStudy (anonymous):

lol thanks. this is just really confusing for me, can you please do it step by step?

OpenStudy (gorv):

@zee_kiya

OpenStudy (gorv):

remember when two number in multiplication have same power ...than can be multiplied directly and power will be same

OpenStudy (gorv):

steps are shown above in start check that ..they are the only steps

OpenStudy (gorv):

got me ??

OpenStudy (anonymous):

no :/

OpenStudy (gorv):

\[\sqrt{x}*\sqrt{y}=\sqrt{x*y}\]

OpenStudy (gorv):

now ??

OpenStudy (anonymous):

yes lol

OpenStudy (gorv):

ok..thatzz zit

OpenStudy (anonymous):

but doesn't it have to be simplified all the way?

OpenStudy (anonymous):